How many moles of H3O+ or OH− must you add to 67.9 mL of strong base solution to adjust its pH from 9.950 to 7.600? Assume a negligible volume change.
1
Expert's answer
2019-07-01T03:55:18-0400

1) Due to the fact that thе pH value decreases (shifts to the acidic region), it is necessary to add Н3О+ ions

2)

"pH = -lgC(\u041d_3\u041e^+)"

сonsequently when pH = 9.950:

с(H3O+) = 10-lgC(H3O+) = 10-9.550 = 1.122*10-10 (mol/L)

When pH = 7.600:

c(H3O+) = 10-7.600 = 2.512*10-8 (mol/L)

3) Now let's calculate the difference in the number of ions H3O+ that must be added to the solution:

Δс(H3O+) = 2.512*10-8 - 1.122*10-10 = 2.501*10-8 (mol/L)

4) If the volume of solution 67.9 mL (0.0679 L):


"n = cV"

n = 2.501*10-8*6.79*10-2 = 16.982*10-10 = 1.698*10-9 (mol)


Answer: 1.698*10-9 mol